OpenStudy (anonymous):

Solve for y: y + 2 = 8 y = 11 y = -11 y = -9 y = 9

OpenStudy (anonymous):

y = 8-2 y = 6 so either your problem is incorrect or the solutions

OpenStudy (anonymous):

2/3y is what got cut off

OpenStudy (anonymous):

Thats right

Directrix (directrix):

y + 2 = 8 y + 2 -2 = 8 - 2 y = 6 Answer not shown as an option

OpenStudy (anonymous):

Solve for y: 2/3y + 2 = 8

OpenStudy (anonymous):